What is the normality of 25.0 g of Al(OH)3 dissolved in 100.0 mL of water?


1
Expert's answer
2021-05-17T04:16:16-0400

"Molarity= \dfrac{\dfrac{25}{78}}{0.1}"


= 3.20M

Hence, Normality= "n\times Molarity"

"3\times 3.20 = 9.60"